# IW-Bench: Evaluating Large Multimodal Models for Converting Image-to-Web

### v1

@@ Abstract
Recently advancements in large multimodal models have led to significant strides in image comprehension capabilities. Despite these advancements, there is a lack of the robust benchmark specifically for assessing the Image-to-Web conversion proficiency of these large models. Primarily, it is essential to ensure the integrity of the web elements generated. These elements comprise visible and invisible categories. Previous evaluation methods (e.g., BLEU) are notably susceptible to significant alterations due to the presence of invisible elements in Web. Furthermore, it is crucial to measure the layout information of web pages, referring to the positional relationships between elements, which is overlooked by previous work. To address challenges, we have curated and aligned a benchmark of images and corresponding web codes (IW-Bench). Specifically, we propose the Element Accuracy, which tests the completeness of the elements by parsing the Document Object Model (DOM) tree. Layout Accuracy is also proposed to analyze the positional relationships of elements by converting DOM tree into a common subsequence. Besides, we design a five-hop multimodal Chain-of-Thought Prompting for better performance, which contains five hop: 1) SoM prompt injection. 2) Inferring Elements. 3) Inferring Layout. 4) Inferring Web code. 5) Reflection. Our benchmark comprises 1200 pairs of images and web codes with varying levels of difficulty. We have conducted extensive experiments on existing large multimodal models, offering insights into their performance and areas for improvement in image-to-web domain.
@@
Hongcheng Guo  Wei Zhang  Junhao Chen  Yaonan Gu Jian YangJunjia DuBinyuan HuiTianyu Liu Jianxin MaChang ZhouZhoujun Li Beihang University  Alibaba Group  Tsinghua University Nanyang Technological University  National University of Singapore
@@ 1 Introduction
The development of large multimodal models has emerged as a new trend, starting with GPT-4 OpenAI (2023b). An increasing number of multimodal models have been introduced Anil et al. (2023); Liu et al. (2023a); Zhu et al. (2023a); Bai et al. (2023); Dai et al. (2023), extending the powerful comprehension capabilities of large language models to multiple tasks Lu et al. (2018); Yang et al. (2023); Sidorov et al. (2020); Guo et al. (2022); Yang et al. (2024). Recently, the task of converting images into web code Laurençon et al. (2024); Patil et al. (2020); Beltramelli (2017) has garnered significant attention due to its impressive performance. In Figure 1, this task tests the synthesis abilities of large multimodal models, encompassing the fine-grained recognition of components within images, the assessment of the relative positioning of elements in webpages, and the capability to generate code. The prowess of large models in generating front-end code has been a source of astonishment, yet, notably, there has been almost no evaluation benchmark related to large multimodal models for this domain.
@@
Constructing a benchmark for appraising Image-to-Web tasks is markedly more intricate than the conventional programming challenges Davody et al. . This heightens complexity stems from the multifaceted nature of web code, encompassing HTML, CSS, and JavaScript. It demands scrutiny not only of functional correctness but also of visual elements like layout, design, and user interaction. Predominant evaluation frameworks for the Image-to-Code task Patil et al. (2020); Bhambure et al. often hinge on similarity metrics such as BLEU. Nevertheless, these metrics encounter substantive limitations in the context of web evaluation. The intricate web of JavaScript functionalities and CSS styling options means that disparate web codes can produce indistinguishable visual outcomes. Consequently, such conventional metrics falter in precisely assessing the completeness of elements depicted in images and their layout intricacies. A further complication arises from non-visible elements, introducing a fragility in these evaluation methods. Varied development ways can lead to substantial discrepancies in these non-visible components, even among web pages that appear visually identical. Recent WebSight Laurençon et al. (2024) is finetuned on numerous image-code pairs, however it has not conducted performance evaluations. These challenges underscore the imperative for a more sophisticated and comprehensive benchmark.
@@ 2 Related work
Multimodal benchmarks for large multimodal models (Bitton et al., 2023; Yu et al., 2023; Liu et al., 2023c; Xu et al., 2023b; Shao et al., 2023) have assess the instruction-following and reasoning capabilities. As these foundation models become more relevant to real-world applications, unlike prior work, we propose IW-Bench to benchmark their generation capabilities of the hot Image-to-Web area on a diverse set of visual contexts. To the best of our knowledge, WebSight Laurençon et al. (2024) is the only relevant benchmark. However, it is more like a fine-tuning dataset for large multimodal models rather than a benchmark for evaluation. The detailed comparison is in Table 1. More related work on image-to-web, large language models, and chain-of-thought methods are in Appendix C.
@@ 3 IW-Bench
We design a pipeline for the construction of IW-Bench in Figure 3. Specifically, to obtain complex-level data, we crawl publicly available web pages and perform de-identification and simplification. For simple and medium level data, we prompt GPT-4 OpenAI (2023a) to generate them. Finally, we filter out some low-quality samples from the obtained image-code pairs. Quantity of IW-Bench is in Table 2 and word cloud is in Figure 2. We present examples in Appendix I.
@@ 3.1 Data Collection
To better quantify differences among models, we opt to create simpler data by prompting GPT4 OpenAI (2023a) following the previous work Wang et al. (2023c); Xu et al. (2023a). The process begins with GPT-4 randomly selecting a relevant domain in Appendix E.1. Based on the desired complexity, GPT-4 then chooses specific elements. Finally, it selects the JavaScript events in Apppendix D.1 to be integrated into the web page. In IW-Bench, 18% of the data comes from real web pages, covering a wide range of types, including public repositories, open-source projects, and websites representing various content categories. The data collection process strictly follows the robots.txt protocol and complies with website policies. To extract images from web code, we use Playwright  [^1] https://github.com/microsoft/playwright.
@@ 3.3 Expert Review
Establishing IW-Bench requires a significant amount of manpower. In the initial collection phase, we hire 20 web engineers to assess complexity levels based on criteria such as HTML tag variety, DOM tree depth, and the presence of scripting. To ensure accuracy, each data point is reviewed by at least three engineers using cross-validation and a majority-rule approach. To address disagreements and maintain high-quality data, we implement several strategies. First, we provide comprehensive training sessions and detailed guidelines to ensure consistency with the criteria. Additionally, we design a consensus mechanism for resolving ties or disagreements. Following the majority-rule principle, engineers vote on the results, and any examples flagged as low quality are re-evaluated until a consensus is reached.
@@ 3.4 Metric Design
We propose Element Accuracy and Layout Accuracy. To effectively compare visible and invisible elements, the initial step involves constructing a Document Object Model (DOM) tree. After the DOM tree is constructed, we employ a consistent traversal strategy to inspect and analyze these elements. During the traversal, we identify each element and collect detailed information about them, such as their tag types, classes, IDs, style attributes, and their hierarchical relationships. This allows us to comprehensively compare visible elements (like text, images, and buttons) with invisible elements (such as script tags and metadata).
For element accuracy, we gain the average score from six perspectives: Tag, Text Content, Attribute, Style, JavaScript, Children.. Specifically, when we compare tag, we only check if the tag names of two elements are identical. If the tag names are an exact match (e.g., both are $<div>$ or $<p>$), the tag_score is 1, otherwise, the tag_score is 0. For the score of text_content, we calculate the content similarity of two elements with SequenceMatcher  [^2] https://github.com/m-matelski/mdiff. To gain the attribute_score, we first define how different HTML tags should compare their attributes, if the corresponding attributes of two elements are an exact match, they score higher.
For style_score, we filter out key style properties (e.g., color, font-size) and ignore default values. Then we compare these filtered style properties, considering them a match if both elements have the same value for the same property. More details of style are in Appendix H. After that, JavaScript event bindings on two elements are compared, we consider specific events (e.g., onclick, onload) and checks if each event is bound to both elements compared to ground truth. The total events are listed in Appendix E. The final score is the number of matching events divided by the total number of events. Let $S_{i}$ be the score from perspective $i$, for $i$in {Tag, Text Content, Attribute, Style, JavaScript, Children} . The set of test elements is $E_{test}$, the set of label elements is $E_{label}$ For each element $E_{j}$ in the testset $E_{test}$, the average score calculated as:
$E_{j}=\frac{1}{n}\sum_{i=1}^{n}S_{i}$ (1)
Given the threshold $T$. The final element accuracy, $EA$, is then:
@@
For layout accuracy, we apply the Longest Common Subsequence (LCS) to derive a web elements into a structured list. Subsequently,list that encapsulates the consistent elements across the layouts. To quantify the similarity index, we divide the length of this common subsequence by the total length of the element list.
$LA=\frac{\text{LCS}(L1,L2)}{\text{Len}(L1)}$ (3)
where $LA$ is the layout accuracy, let $L1$ be the list of web elements from the label HTML, and $L2$ be the list from the generated Web code.
@@ 4.1 SoM Prompt Injection
Motivated by previous work Yang et al. (2023), we aim to enhance images with Scenes of Meaning (SoM) as shown in Figure 5. We first identify key elements to understand the core theme. Using specialized tools, we seamlessly integrate SoM cues like text labels, arrows, and highlights into the images. These cues guide multimodal models to focus on important elements, improving image comprehension and precision in recognizing fine details.
@@ 4.2 Inferring Elements
After integrating the SoM prompts, we can more precisely infer the elements within the image. This enhanced inference covers various aspects, including categorizing elements based on their types and understanding their functional roles within the image. SoM prompts provide significant assistance, enabling us to delve deeper into understanding both visible and invisible elements in the image, as well as their contributions to the overall meaning.
@@ 4.3 Inferring Layout
In the third hop, we prompt the models to infer the layout information in the image. This includes a detailed analysis of the spatial arrangement and organization of the web elements depicted in the image. We examine factors such as the alignment, proximity, and hierarchy of elements to understand their functional relationships and visual impact. This phase is crucial for understanding the overall structure and flow of the user interface, enabling models to recognize how users might interact with and navigate through the web content. By accurately decoding the layout information, large multimodal models can gain insights into the applied design principles and subsequently refine their generated content.
@@ 5.1 Experiment setting
We select recent multimodal language models as baselines. GPT4V (gpt-4-1106-preview) (OpenAI, 2023b), Qwen-VL-Chat,Qwen-VL-Plus, Qwen-VL-Max Bai et al. (2023), miniGPT-4-LLaMA-2-7B Zhu et al. (2023a), LLaVA-LLaMA-2-13B Liu et al. (2023a), mPLUG-OWL2 Ye et al. (2023), LLaMA-Adapter-V2-7B Gao et al. (2023), WebSight Laurençon et al. (2024), Gemini Pro Anil et al. (2023), Claude3 Opus  [^3] https://www.anthropic.com/news/claude-3-family. All experiments are conducted with 16 NVIDIA A100 GPUs (80G). The $T$ in our experiment is 0.9. $N$ is default set to be 3 and we choose the best result. We do not conduct Five-hop MCoT on Websight model as it fails to accommodate simultaneous input of images and text.
@@ 5.2 Experimental Results
Our results in Table 5 are divided into two main sections: the first section outlines the results of web code directly generated by the model while the second section details outcomes following the five-hop multi-modal chain-of-thought technique. We evaluate the efficacy of various models by examining their accuracy in identifying elements and arranging layouts across three distinct levels of complexity: simple, medium, and complex.
@@ 6.2 Ablation on Reflection
Figure 7: Visualization on the reflection. The reflection times $N$=3. We can clearly see the element and layout improvements during different times of reflection.
@@ Appendix A Data De-identification
Appendix B Limitations
Language Scope.
Data Quantity.
@@ C.1 Image-to-Web
Recent advancements in the field of web development have seen researchers exploring innovative ways to convert webpage images into HTML code Patil et al. (2020); Bhambure et al. ; Davody et al. . Researcher Patil et al. (2020) focuses on automatically generating HTML code from webpage mock-ups. This involves the use of pix2code Beltramelli (2017), which trains a CNN-based model on various webpage structure mock-ups, showcasing the potential of automating GUI creation. Another method Aşıroğlu et al. (2019) employs computer vision to identify objects and deep systematic analysis for result generation. Sketch2Code Jain et al. (2019) divides the problem into three parts: object recognition, bounding box creation, and the creation of a functional prototype application. Recently, attention has shifted towards GPT4V for its image understanding and code generation capabilities. In the project screenshot-to-code  [^4] https://github.com/abi/screenshot-to-code, GPT4V OpenAI (2023b) is used as the base model. WebSight Laurençon et al. (2024) from Huggingface is finetuned on numerous collected image-code pairs, however it has not conducted public performance evaluations. We desperately need a benchmark to evaluate large multimodal models in this domain.
C.2 Large Language Model
Recent advancements in AI have led to the development of generative foundation models (Bommasani et al., 2021) like GPT-3, ChatGPT, GPT-4, Claude and LLaMA (Brown et al., 2020; OpenAI, 2022, 2023a; Anthropic, 2023; Touvron et al., 2023; Zhang et al., 2023b), which excel in a variety of text-based tasks without specific finetuning. Their performance has been evaluated across disciplines such as QA, math, and science (Chen et al., 2021; Sun et al., 2023; Wang et al., 2023b; Huang et al., 2023, 2022; Liu et al., 2023b; Guo et al., 2024; Zhang et al., 2023a). On the vision-language side, there are several generative foundation models such as Qwen-Max, Qwen-VL, LLaVA, MiniGPT4, InstructBLIP, Flamingo,, Multimodal Bard (Bai et al., 2023; Zhu et al., 2023a; Dai et al., 2023; Alayrac et al., 2022; Awadalla et al., 2023; Gao et al., 2023; Google, 2023) that are trained on extensive image-text data, paving the way for multimodal learning (Schuhmann et al., 2022; Sharma et al., 2018; Lin et al., 2014; Zhu et al., 2023b). In addition, models specialized versions for document understanding are proposed (Zhang et al., 2023c; Ye et al., 2023). Benchmarks, like Visit-Bench, MMBench (Bitton et al., 2023; Yu et al., 2023; Liu et al., 2023c; Xu et al., 2023b; Shao et al., 2023), have assess the instruction-following and reasoning capabilities. As these foundation models become more relevant to real-world applications, unlike prior work, we plan to benchmark their capabilities of the hot Image-to-Web area on a diverse set of visual contexts.

### v2

@@ Abstract
Recently advancements in large multimodal models have led to significant strides in image comprehension capabilities. Despite these advancements, there is a lack of the robust benchmark specifically for assessing the Image-to-Web conversion proficiency of these large models. Primarily, it is essential to ensure the integrity of the web elements generated. These elements comprise visible and invisible categories. Previous evaluation methods (e.g., BLEU) are notably susceptible to significant alterations due to the presence of invisible elements in Web. Furthermore, it is crucial to measure the layout information of web pages, referring to the positional relationships between elements, which is overlooked by previous work. To address challenges, we have curated and aligned a benchmark of images and corresponding web codes (IW-Bench). Specifically, we propose the Element Accuracy, which tests the completeness of the elements by parsing the Document Object Model (DOM) tree. Layout Accuracy is also proposed to analyze the positional relationships of elements by converting DOM tree into a common subsequence. Besides, we design a five-hop multimodal Chain-of-Thought Prompting for better performance, which contains five hop: 1) SoM prompt injection. 2) Inferring Elements. 3) Inferring Layout. 4) Inferring Web code. 5) Reflection. Our benchmark comprises 1200 pairs of images and web codes with varying levels of difficulty. We have conducted extensive experiments on existing large multimodal models, offering insights into their performance and areas for improvement in image-to-web domain.  [^1] We provide code and dataset at https://github.com/HC-Guo/IWBench
@@
Hongcheng Guo1, Wei Zhang1, Junhao Chen3, Yaonan Gu5, Jian Yang2, Junjia Du4, Shaosheng Cao, Binyuan Hui2, Tianyu Liu2, Jianxin Ma2, Chang Zhou2, Zhoujun Li1thanks: Corresponding author., 1Beihang University 2Alibaba Group 3Tsinghua University 4Nanyang Technological University 5National University of Singapore {lizj, hongchengguo}@buaa.edu.cn,
@@ 1 Introduction
The development of large multimodal models has emerged as a new trend, starting with GPT-4 openai2023gpt4v. An increasing number of multimodal models have been introduced gemini; liu2023llava; zhu2023minigpt; qwenvl; instructblip, extending the powerful comprehension capabilities of large language models to multiple tasks lu2018rvqa; som; sidorov2020textcaps. Recently, the task of converting images into web code websight; imagecode1; pix2code has garnered significant attention due to its impressive performance. In Figure 1, this task tests the synthesis abilities of large multimodal models, encompassing the fine-grained recognition of components within images, the assessment of the relative positioning of elements in webpages, and the capability to generate code. The prowess of large models in generating front-end code has been a source of astonishment, yet, notably, there has been almost no evaluation benchmark related to large multimodal models for this domain.
@@
Constructing a benchmark for appraising Image-to-Web tasks is markedly more intricate than the conventional programming challenges imagecode7. This heightens complexity stems from the multifaceted nature of web code, encompassing HTML, CSS, and JavaScript. It demands scrutiny not only of functional correctness but also of visual elements like layout, design, and user interaction. Predominant evaluation frameworks for the Image-to-Code task imagecode1; imagecode3 often hinge on similarity metrics such as BLEU. Nevertheless, these metrics encounter substantive limitations in the context of web evaluation. The intricate web of JavaScript functionalities and CSS styling options means that disparate web codes can produce indistinguishable visual outcomes. Consequently, such conventional metrics falter in precisely assessing the completeness of elements depicted in images and their layout intricacies. A further complication arises from non-visible elements, introducing a fragility in these evaluation methods. Varied development ways can lead to substantial discrepancies in these non-visible components, even among web pages that appear visually identical. Recent WebSight websight is finetuned on numerous image-code pairs, however it has not conducted performance evaluations. These challenges underscore the imperative for a more sophisticated and comprehensive benchmark.
@@ 2 Related work
Multimodal benchmarks for large multimodal models (bitton2023visit; yu2023mm; liu2023mmbench; xu2023lvlm; shao2023tiny) have assess the instruction-following and reasoning capabilities. As these foundation models become more relevant to real-world applications, unlike prior work, we propose IW-Bench to benchmark their generation capabilities of the hot Image-to-Web area on a diverse set of visual contexts. To the best of our knowledge, WebSight websight is the only relevant benchmark. However, it is more like a fine-tuning dataset for large multimodal models rather than a benchmark for evaluation. The detailed comparison is in Table 1. More related work on image-to-web, large multimodal models, and chain-of-thought methods are in Appendix C.
@@ 3 IW-Bench
We design a pipeline for the construction of IW-Bench in Figure 3. Specifically, to obtain complex-level data, we crawl publicly available web pages and perform de-identification and simplification. For simple and medium level data, we prompt GPT-4 openai2023gpt4 to generate them. Finally, we filter out some low-quality samples from the obtained image-code pairs. Quantity of IW-Bench is in Table 2 and word cloud is in Figure 2. We present examples in Appendix I.
@@ 3.1 Data Collection
To better quantify differences among models, we opt to create simpler data by prompting GPT4 openai2023gpt4 following the previous work self_instruct; xu2023baize. The process begins with GPT-4 randomly selecting a relevant domain in Appendix E.1. Based on the desired complexity, GPT-4 then chooses specific elements. Finally, it selects the JavaScript events in Apppendix D.1 to be integrated into the web page. In IW-Bench, 18% of the data comes from real web pages, covering a wide range of types, including public repositories, open-source projects, and websites representing various content categories. The data collection process strictly follows the robots.txt protocol and complies with website policies. To extract images from web code, we use Playwright  [^2] https://github.com/microsoft/playwright.
@@ 3.3 Expert Review
Establishing IW-Bench involves substantial manpower. In the collection phase, 20 web engineers assess complexity based on HTML tag variety, DOM tree depth, and scripting presence. Each data point is cross-validated by three engineers, with disagreements resolved through training, detailed guidelines, and a majority-rule consensus mechanism. Low-quality examples are re-evaluated until agreement is achieved.
@@ 3.4 Metric Design
To evaluate the fidelity of web elements, we propose two novel metrics: Element Accuracy (EA) and Layout Accuracy (LA), which are designed to compare visible and invisible elements in web pages. The evaluation process begins with constructing a Document Object Model (DOM) tree for the test and reference web pages, followed by a structured traversal and analysis of the elements. The DOM tree traversal systematically inspects each element, collecting detailed information such as tag types, attributes, styles, JavaScript bindings, and hierarchical relationships. This unified approach enables the evaluation of visible elements (e.g., text, images, buttons) alongside invisible ones (e.g., <script> tags, metadata).
3.4.1 Element Accuracy (EA)
Element Accuracy measures the similarity between corresponding elements in the test and reference web pages based on six key attributes: Tag, Text Content, Attributes, Style, JavaScript, and Children. Let the test element set be $E_{\text{test}}$, and the label element set be $E_{\text{label}}$. For each test element $E_{j}$, an average score $S_{j}$ is computed across these six perspectives. The Element Accuracy is defined as:
@@ 3.4.1 Element Accuracy (EA)
where $T$ is the threshold, and $\mathbb{1}(E_{j}>T)$ is an indicator function that equals 1 if $E_{j}$ exceeds $T$, and 0 otherwise.
The similarity scores for each perspective are computed as follows:
  • Tag: Tags are compared by their names:
    $S_{\text{tag}}=\begin{cases}1&\text{if }\text{tag}_{\text{test}}=\text{tag}_{\text{label}},\\ 0&\text{otherwise}.\end{cases}$ (2)
  • Text Content: The similarity of text content between two elements is computed by SequenceMatcher (SM)  [^3] https://github.com/m-matelski/mdiff:
    $S_{\text{text}}=\text{SM}(\text{content}_{\text{test}},\text{content}_{\text{label}})$ (3)
  • Attributes: A predefined mapping is used for each tag type. If the corresponding attributes of two elements match exactly, they receive a higher score:
    $S_{\text{attr}}=\frac{\sum_{k\in A}\mathbb{1}(\text{attr}_{k}^{\text{test}}=\text{attr}_{k}^{\text{label}})}{|A|}$ (4)
    where $A$ is the set of relevant attributes for the tag.
  • Style: Key style properties (e.g., color, font-size) are filtered out, ignoring default values. Matching values contribute to the score:
    $S_{\text{style}}=\frac{\sum_{p\in P}\mathbb{1}(\text{style}_{p}^{\text{test}}=\text{style}_{p}^{\text{label}})}{|P|}$ (5)
    where $P$ is the set of key style properties.
  • JavaScript: Event bindings on two elements are compared (e.g., onclick, onload). Let $J$ be the set of relevant events. The JavaScript score is:
    $S_{\text{js}}=\frac{\sum_{e\in J}\mathbb{1}(\text{event}_{e}^{\text{test}}=\text{event}_{e}^{\text{label}})}{|J|}$ (6)
  • Children: The hierarchical structure is evaluated using a tree-edit distance algorithm.
The average score for each test element $E_{j}$ is:
$E_{j}=\frac{\sum_{i=1}^{n}S_{i}}{n},$ (7)
where $n=6$ represents the number of evaluated attributes: Tag, Text Content, Attribute, Style, JavaScript, Children.
3.4.2 Layout Accuracy (LA)
Layout Accuracy quantifies the structural similarity between the layouts of web pages by applying the Longest Common Subsequence (LCS) on their element lists. The layout is linearized into a sequence of elements for comparison.
Let $L_{1}$ be the list of web elements from the label HTML, and $L_{2}$ be the list from the generated Web code. The layout accuracy is defined as:
$LA=\frac{\text{LCS}(L_{1},L_{2})}{\text{Len}(L_{1})}$ (8)
$\text{LCS}(L_{1},L_{2})=\max\left(\delta(e_{1},e_{2})+\text{LCS}(L_{1}^{\prime},L_{2}^{\prime})\right)$ (9)
where $\text{LCS}(L_{1},L_{2})$ is the length of the longest common subsequence of elements between $L_{1}$ and $L_{2}$, $\text{Len}(L_{1})$ is the total length of $L_{1}$. $\delta(e_{1},e_{2})$ be a similarity function between two elements $e_{1}$ and $e_{2}$. $L_{1}^{\prime}$ and $L_{2}^{\prime}$ are the remaining subsequences after matching $e_{1}$ and $e_{2}$.
Further details about style properties are in Appendix H, and JavaScript events are in Appendix E.
@@ 4.1 SoM Prompt Injection
Inspired by prior work som, we enhance images with Scenes of Meaning (SoM) cues, as shown in Figure 5. By identifying key elements and integrating cues like text labels, arrows, and highlights, we guide multimodal models to focus on crucial details, improving comprehension and precision.
@@ 4.2 Inferring Elements
Integrating SoM prompts enhances element inference in images, improving categorization by type and functional role. This aids in understanding both visible and invisible elements and their contributions to the overall context.
@@ 4.3 Inferring Layout
In the third hop, models infer layout information, analyzing spatial arrangement, alignment, proximity, and hierarchy of web elements. This step uncovers functional relationships and visual impact, helping models understand UI structure, user interaction flow, and design principles to refine content generation.
@@ 5.1 Experiment setting
We select recent multimodal language models as baselines. GPT4V (gpt-4-1106-preview) (openai2023gpt4v), Qwen-VL-Chat,Qwen-VL-Plus, Qwen-VL-Max qwenvl, miniGPT-4-LLaMA-2-7B zhu2023minigpt, LLaVA-LLaMA-2-13B liu2023llava, mPLUG-OWL2 ye2023mplug, LLaMA-Adapter-V2-7B gao2023llamaadapterv2, WebSight websight, Gemini Pro gemini, Claude3 Opus  [^4] https://www.anthropic.com/news/claude-3-family. All experiments are conducted with 16 NVIDIA A100 GPUs (80G). The $T$ in our experiment is 0.9. $N$ is default set to be 3 and we choose the best result. We do not conduct Five-hop MCoT on Websight model as it fails to accommodate simultaneous input of images and text.
@@ 5.2 Experimental Results
Our results in Table 5 are divided into two sections: (1) web code directly generated by the model, and (2) results using the five-hop multi-modal chain-of-thought. We evaluate model in identifying elements and arranging layouts across three complexity levels: simple, medium, and complex.
@@ 6.2 Ablation on Reflection
Figure 7: Visualization on the reflection. The reflection times NN=3. We can clearly see the element and layout improvements during different times of reflection.
@@ Appendix A Data De-identification
Appendix B Crowdsourcing
In our study, we recognize several potential risks for participants. First, there is a risk to privacy and confidentiality since participants are required to share personal information. To address this, we anonymize all data and store it securely, ensuring access is limited to authorized personnel only. Second, psychological risks, such as discomfort or stress, may arise during the tasks. We mitigate these by providing clear instructions and holding debriefing sessions to support participants throughout the study. Additionally, participants are free to withdraw at any time without any penalties. While no significant physical risks are associated with our procedures, we continuously monitor for any signs of distress and offer appropriate support as needed. Each participant receives hourly compensation of \$20. Our primary participants for this study are web engineers.
@@ C.1 Image-to-Web
Recent advancements in the field of web development have seen researchers exploring innovative ways to convert webpage images into HTML code imagecode1; imagecode3; imagecode7. Researcher imagecode1 focuses on automatically generating HTML code from webpage mock-ups. This involves the use of pix2code pix2code, which trains a CNN-based model on various webpage structure mock-ups, showcasing the potential of automating GUI creation. Another method autoimagecode4 employs computer vision to identify objects and deep systematic analysis for result generation. Sketch2Code imagecode6 divides the problem into three parts: object recognition, bounding box creation, and the creation of a functional prototype application. Recently, attention has shifted towards GPT4V for its image understanding and code generation capabilities. In the project screenshot-to-code  [^5] https://github.com/abi/screenshot-to-code, GPT4V openai2023gpt4v is used as the base model. WebSight websight from Huggingface is finetuned on numerous collected image-code pairs, however it has not conducted public performance evaluations. We desperately need a benchmark to evaluate large multimodal models in this domain.
C.2 Multimodal Large Language Model
Recent advancements in AI have led to the development of generative foundation models (bommasani2021opportunities) like GPT-3, ChatGPT, GPT-4, Claude and LLaMA (brown2020language; openai2022chatgpt; openai2023gpt4; claude2; touvron2023llama; llamaadapter2023), which excel in a variety of text-based tasks without specific finetuning. Their performance has been evaluated across disciplines such as QA, math, and science (chen2021evaluating; sun2023scieval; wang2023scibench; huang2023c; huang2022language; liu2023agentbench). On the vision-language side, there are several generative foundation models such as Qwen-Max, Qwen-VL, LLaVA, MiniGPT4, InstructBLIP, Flamingo,, Multimodal Bard (qwenvl; zhu2023minigpt; instructblip; alayrac2022flamingo; awadalla2023openflamingo; gao2023llamaadapterv2; google2023bard) that are trained on extensive image-text data, paving the way for multimodal learning (schuhmann2022laion; sharma2018conceptual; lin2014microsoft; zhu2023multimodal). In addition, models specialized versions for document understanding are proposed (zhang2023llavar; ye2023mplug). Benchmarks, like Visit-Bench, MMBench (bitton2023visit; yu2023mm; liu2023mmbench; xu2023lvlm; shao2023tiny), have assess the instruction-following and reasoning capabilities. As these foundation models become more relevant to real-world applications, unlike prior work, we plan to benchmark their capabilities of the hot Image-to-Web area on a diverse set of visual contexts.